# FACT-GS: Frequency-Aligned Complexity-Aware Texture Reparameterization for 2D Gaussian Splatting

### v2

@@ Abstract
Realistic scene appearance modeling has advanced rapidly with Gaussian Splatting, which enables real-time, high-quality rendering. Recent advances introduced per-primitive textures that incorporate spatial color variations within each Gaussians, improving their expressiveness. However, texture-based Gaussians parameterize appearance with a uniform per-Gaussian sampling grid, allocating equal sampling density regardless of local visual complexity. This leads to inefficient texture space utilization, where high-frequency regions are under-sampled and smooth regions waste capacity, causing blurred appearance and loss of fine structural detail. We introduce FACT-GS, a Frequency-Aligned Complexity-Aware Texture Gaussian Splatting framework that allocates texture sampling density according to local visual frequency. Grounded in adaptive sampling theory, FACT-GS reformulates texture parameterization as a differentiable sampling-density allocation problem, replacing the uniform textures with a learnable frequency-aware allocation strategy implemented via a deformation field whose Jacobian modulates local sampling density. Built on 2D Gaussian Splatting, FACT-GS performs non-uniform sampling on fixed-resolution texture grids, preserving real-time performance while recovering sharper high-frequency details under the same parameter budget.
@@ 1 Introduction
Realistic 3D appearance modeling is essential for large-scale reconstruction, dynamic scene understanding, and photorealistic rendering [kulhanek2024wildgaussians, lin2024vastgaussian, fischer2024dynamic], but achieving high fidelity under real-time constraints remains challenging. Implicit volumetric fields such as NeRF [mildenhall2021nerf, barron2021mip, wu2024dynamic, zhao2023instant] provide high visual quality but require dense sampling and costly MLP inference, limiting scalability. In contrast, 3D Gaussian Splatting (3DGS) [kerbl20233d, yu2024mip, feng2025flashgs] represents scenes using explicit differentiable Gaussian primitives, achieving real-time high-quality rendering. Building on this, 2D Gaussian Splatting [huang20242d] uses planar primitives that simplify ray–primitive interaction and improve geometric stability. However, 2D and 3D Gaussian splatting represent appearance using view-dependent spherical harmonics [kerbl20233d, huang20242d], without spatial color variation within each primitive, constraining its expressiveness.
@@
To address this limitation, texture-based Gaussian extensions [chao2025textured, rong2025gstex, huang2024textured] attach per-Gaussian spatially varying textures. However, the appearance parameterization of these methods is uniformly distributed across the texture space of each Gaussian, without adapting to local visual complexity [hachisuka2008multidimensional, sztrajman2021neural, wang2011efficiency]. This uniform allocation leads to inefficient utilization of the texture space: high-frequency regions (e.g., sharp edges, fine patterns, numerical markings) are allocated insufficient texture space, resulting in detail loss, while large smooth regions inefficiently consume capacity to represent nearly uniform color. In practice, each Gaussian can only store a low-resolution texture patch (e.g., $4\times 4\times 4$), exacerbating the expressiveness constraints imposed by uniform sampling. As shown in Fig. LABEL:fig:teaser, this results in blurred appearances in visually complex regions. Simply increasing texture resolution yields only marginal improvements at quadratically increased memory and bandwidth cost, while neural texture fields [xu2024texture, zhang2025neural] break the real-time rendering performance of the Gaussian-based representations.
@@
Motivated by this structural inefficiency, we propose an adaptive sampling scheme of the per-Gaussian texture where sampling should follow the local signal frequency rather than remain spatially uniform. To achieve this, we introduce a frequency-aligned differentiable texture reparameterization that continuously modulates sampling density across the texture domain. Grounded in adaptive sampling theory [heckbert1989fundamentals, rousselle2011adaptive, zwicker2015recent], this allocates more sampling capacity to high-frequency regions and less to smooth areas. Building on this formulation, we present FACT-GS, a Frequency-Aligned Complexity-Aware Texture Gaussian Splatting framework that integrates this frequency-adaptive reparameterization.
@@
Concretely, FACT-GS is built upon 2D Gaussian Splatting [huang20242d] and introduces a learnable deformation field defined in the texture domain. This field predicts continuous per-texel displacements, inducing a smooth warping whose Jacobian determinant [sander2001texture] directly modulates local sampling density. In contrast to previous uniform texture parameterizations [chao2025textured, rong2025gstex, huang2024textured], which allocate equal resolution regardless of signal complexity, our warping adaptively concentrates samples in high-frequency regions while compressing low-frequency areas. This realizes adaptive sampling as a differentiable, per-primitive reparameterization, allowing texture capacity to naturally focus where visual detail is needed. The entire framework is trained end-to-end under standard photometric supervision, as the deformation field is designed to integrate seamlessly into the differentiable splatting pipeline. As a result, FACT-GS recovers sharper high-frequency appearance and uses parameters more efficiently, all while maintaining real-time rendering performance under the same primitive budget.
@@ 2 Related Work
Novel View Synthesis. Novel view synthesis (NVS) aims to generate photorealistic renderings of 3D scenes from unseen viewpoints, typically given multi-view images with known camera poses. Its development originates from classical SfM and MVS pipelines [snavely2006photo, snavely2008modeling, agarwal2011building, schonberger2016structure, seitz2006comparison, furukawa2009accurate, furukawa2010towards], which reconstruct geometry through multi-view correspondences but suffer from occlusion sensitivity and limited viewpoint extrapolation. Neural radiance fields (NeRF) [mildenhall2021nerf] later revolutionized NVS by representing scenes as implicit neural functions optimized via photometric supervision. Subsequent extensions improve anti-aliasing [barron2021mip], large-scale reconstruction [tancik2022block, barron2022mip], reflection modeling [verbin2022ref], and acceleration via grid-based encodings [zhao2023instant], yet they remain constrained by dense ray sampling and slow MLP inference.
This motivates explicit representations such as 3D Gaussian Splatting [kerbl20233d], which replace neural volumes with differentiable Gaussian primitives for real-time rendering. Building upon this foundation, follow-up works improve geometric accuracy and rendering stability through enhanced kernel formulations and densification strategies [fu2024colmap, lei2025mosca, li20253d]. To further refine geometry, 2D Gaussian Splatting [huang2024textured] introduces planar Gaussian primitives for more accurate and efficient radiance field representation, enabling fine-grained geometric reconstruction while maintaining high appearance quality. However, both 3D and 2D Gaussian Splatting modeled appearance via view-dependent spherical harmonic coefficients, but lack spatial color variation within each primitive, limiting their expressiveness.
Spatially Varying Gaussians. Recent works introduce spatially varying Gaussians to overcome this constraint. One line of research augments each primitive with RGB or RGBA textures to model spatially varying colors [chao2025textured, xu2024supergaussians, rong2025gstex, huang2024textured, song2024hdgs] or surface normals for relighting [sun2025svg]. Another employs neural texture fields to decouple geometry and appearance [xu2024texture, zhang2025neural], achieving high visual fidelity at the cost of slower training and rendering due to MLP inference.
@@
Adaptive Sampling and Frequency-Aware Parameterization. Classical rendering and Monte Carlo integration extensively explored adaptive sampling [hachisuka2008multidimensional, zwicker2015recent, rousselle2011adaptive, sztrajman2021neural, wang2011efficiency], where sample density is modulated by local signal frequency or reconstruction error to achieve efficient high-frequency reconstruction. At the representation level, multi-resolution texture filtering such as mipmapping and anisotropic filtering [heckbert2007survey, williams1983pyramidal, pharr2024filtering] pre-filters textures across scales to suppress aliasing and maintain visual consistency under varying viewing distances. Recent neural and Gaussian-based formulations [li2024dngaussian, zhu2024fsgs, liu20243dgs] extend this principle to learned 3D representations, enforcing frequency- or scale-consistent parameterization in both geometry and appearance reconstruction. However, these methods operate at the global or feature-map level and still assume uniform sampling within each primitive, limiting their ability to adapt texture-space resolution to local signal variation.
@@ 3 Preliminaries
We adopt the 2D Gaussian Splatting representation [huang20242d]. Each primitive lies on a local tangent plane with center $p_{k}$, orthonormal basis $(t_{\beta},t_{\gamma},n_{k})$, and scales $(s_{\beta},s_{\gamma})$. Local coordinates on the plane are $(\beta,\gamma)$:
@@ 3.2 Textured Gaussians
Standard 2D Gaussians provide only a single view-dependent color for each primitive. Textured Gaussians [chao2025textured] instead attach a learnable RGBA texture map $\mathbf{T}_{i}\in\mathbb{R}^{\tau\times\tau\times 4}$ to each primitive, enabling spatially varying appearance within tangent plane. Although originally formulated in 3D, the official implementation adopts a 2D variant for more accurate and efficient ray–Gaussian intersection, which we also follow.
@@ 4.1 Motivation
Let $C(u,v)$ denote the ground-truth color function, and let the texture $\mathbf{T}$ be its discrete representation, where texels sample $C(u,v)$ in texture space. The sampling density $\rho(u,v)$ represents the number of texels per unit area. Uniform parameterization is structurally inefficient because it assigns the same density everywhere despite large spatial variation in the underlying appearance signal [sander2001texture, zwicker2015recent].
@@
Following prior work that uses local variation as a proxy for frequency content [eskicioglu2002image, rousselle2011adaptive], we adopt the gradient-based target density
@@ 4.2 Frequency-Aligned Complexity-Aware Texture Reparameterization
In contrast to texture-based Gaussians [chao2025textured, rong2025gstex], which operate on fixed uniform grids and primarily improve detail by increasing resolution, our formulation adjusts sampling density geometrically through $\Phi$ while keeping the grid resolution fixed. This enables non-uniform, complexity-aware texture parameterization that remains fully differentiable and end-to-end trainable.
@@
The local signal frequency can be approximated by the gradient magnitude $\|\nabla c_{i}^{\mathrm{tex}}(u,v)\|$ [eskicioglu2002image]. Under the warp, the uniform-grid texture gradient $\nabla\mathbf{T}_{i}^{\mathrm{RGB}}$ is transformed by the singular values of $J_{\Phi_{i}}$, which respectively amplify or attenuate its magnitude. These effects manifest as regions receiving more texels allocated or fewer texels received, enabling the FACT texture $c_{i}^{\mathrm{tex}}(u,v)$ to respond faithfully to higher-frequency variation in the ground-truth color function $C(u,v)$.
@@ 6   Sample $c_{i}^{\mathrm{tex}},\alpha_{i}^{\mathrm{tex}}$ from $\mathbf{T}_{i}(u^{\prime},v^{\prime})$;
Figure 3: Novel view synthesis comparison under reduced primitive budgets (10% / 1% of default 2DGS). 2DGS [huang20242d] loses high-frequency appearance, and Textured GS [chao2025textured] blurs fine detail due to uniform texture sampling. Our method preserves sharp edges, textures, and material details (zoom-ins), achieving significantly higher perceptual fidelity without increasing parameters.
@@ 4.3 Optimization
Following Textured Gaussians [chao2025textured], we adopt a two-stage optimization for stability. In the first stage, the 2D Gaussian primitives are trained using the standard losses from 3D Gaussian Splatting [kerbl20233d], combining $\mathcal{L}_{1}$ photometric and structural similarity ($\mathcal{L}_{\text{SSIM}}$) terms, augmented with a mask loss $\mathcal{L}_{\alpha}$ when ground-truth masks are available:
@@ 5.1 Implementation Details
We build on the gsplat [ye2025gsplat] framework and extend it with a lightweight CUDA kernel for frequency-aligned texture warping. The per-Gaussian texture space sampling-density modulation described in section 4.2 introduces negligible overhead during inference. All experiments are run on a single NVIDIA RTX A6000 GPU.
@@ 5.2 Datasets and Evaluation Metrics
Dataset. We evaluate our method on five standard benchmarks for novel view synthesis: NeRF Synthetic [mildenhall2021nerf] (8 scenes), MipNeRF 360 v2 [barron2022mip] (7 scenes), DTU [jensen2014large] (4 scenes), Tanks & Temples [knapitsch2017tanks] (3 scenes), and LLFF [mildenhall2019llff] (2 scenes).
Metrics. Following Textured GS [chao2025textured], we report Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS) [zhang2018unreasonable] for quantitative evaluation.
@@ 5.3 Experimental Results
Methods NeRF Synthetic [mildenhall2021nerf] MipNeRF 360v2 [barron2022mip] DTU [jensen2014large] Tanks & Temples [knapitsch2017tanks] LLFF [mildenhall2019llff]
@@
Table 1: Quantitative comparison with 2DGS [huang20242d] and Textured GS [chao2025textured] under varying numbers of Gaussian primitives. Our method consistently achieves higher fidelity across different primitive budgets. (Best scores are shown in bold, second-best are underlined.)
Quantitative Results. We evaluate our method against Textured GS [chao2025textured] and 2D GS [huang20242d] under varying primitive budgets ($0\%, 1010\%$, $1\%$). All methods follow the same 2D GS training protocol [huang20242d]. Table 1 summarizes the results; corresponding average primitive counts are provided in the appendix. To ensure comparable parameter budgets, we match per-Gaussian texture capacity by setting $\tau_{\text{tex}}=5$ for Textured GS and $\tau_{\text{FACT}}=4$ for our method.
@@
We further evaluate our method ($100\%$ primitive budget) on the NeRF Synthetic [mildenhall2021nerf] and MipNeRF 360-Indoor [barron2022mip] datasets. We compare against Textured GS [chao2025textured], SuperGS [xu2024supergaussians], GsTex [rong2025gstex], and Neural Shell [zhang2025neural].
As shown in Table 2, our method achieves the highest overall fidelity while maintaining real-time rendering performance (in Table 3). In contrast, methods such as Neural Shell [zhang2025neural] improve accuracy only by increasing computation. This indicates that our method does not rely on capacity scaling, but instead reallocates texture resolution to high-frequency regions—yielding structural efficiency gains directly predicted by adaptive sampling theory.
Detail Preservation Under Limited Primitive Budgets. Fig. 3 compares novel view synthesis when using only $10\%$ and $1\%$ of the default 2DGS primitive count. 2D GS [huang20242d] loses fine textures entirely due to its single-color appearance model, and Textured GS [chao2025textured] produces blurred patterns because its uniform texture grid allocates capacity evenly across the surface. In contrast, our method preserves sharp edges, printed characters, and high-frequency materials by reallocating sampling density to regions that require detail. This confirms that frequency-aligned parameterization maintains perceptual fidelity even when representational resources are severely constrained.
@@
Methods NeRF Synthetic [mildenhall2021nerf] MipNeRF 360-indoor [barron2022mip]
@@
3DGS [kerbl20233d] 33.34 0.969 0.030 31.03 0.921 0.188
SuperGS [xu2024supergaussians] 33.71 0.970 0.031 30.23 0.917 0.188
GsTex [rong2025gstex] 33.37 0.965 0.041 30.46 0.915 0.204
Neural Shell [zhang2025neural] 33.50 0.967 0.032 30.59 0.911 0.174
Textured GS [chao2025textured] (100%) 33.91 0.968 0.024 31.71 0.938 0.061
@@
Table 2: Quantitative comparison with prior Gaussian-based methods, including 3DGS [kerbl20233d], Neural Shell [zhang2025neural], SuperGS [xu2024supergaussians], GsTex [rong2025gstex], and Textured GS [chao2025textured]. All methods are evaluated using the full (100%) 2DGS primitive budget [huang20242d] on NeRF Synthetic [mildenhall2021nerf] and MipNeRF 360-Indoor [barron2022mip]. (Best results are shown in bold, and second-best are underlined.)
Figure 4: Qualitative comparison under high primitive budgets ($100\%$ for Lego, $50\%$ for Church). 2D GS [huang20242d] lacks spatial texture variation, and Textured GS [chao2025textured] shows blurred high-frequency patterns due to uniform texture sampling. Our method preserves thin structures and fine texture details.
Inference Speed (FPS) Textured GS [huang2024textured] Neural Shell [zhang2025neural] Ours

### v3

@@ Abstract
Realistic scene appearance modeling has advanced rapidly with Gaussian Splatting, which enables real-time, high-quality rendering. Recent advances introduced per-primitive textures that incorporate spatial color variations within each Gaussian, improving their expressiveness. However, texture-based Gaussians parameterize appearance with a uniform per-Gaussian sampling grid, allocating equal sampling density regardless of local visual complexity, which leads to inefficient texture space utilization. We introduce FACT-GS, a Frequency-Aligned Complexity-Aware Texture Gaussian Splatting framework that allocates texture sampling density according to local visual frequency. Grounded in adaptive sampling theory, FACT-GS reformulates texture parameterization as a differentiable sampling-density allocation problem, replacing the uniform textures with a learnable frequency-aware allocation strategy implemented via a deformation field whose Jacobian modulates local sampling density. Built on 2D Gaussian Splatting, FACT-GS performs non-uniform sampling on fixed-resolution texture grids, preserving real-time performance while recovering sharper high-frequency details under the same parameter budget.
@@ 1 Introduction
Realistic 3D appearance modeling is essential for large-scale reconstruction, dynamic scene understanding, and photorealistic rendering [19, 23, 7], but achieving high fidelity under real-time constraints remains challenging. Implicit volumetric fields such as NeRF [25, 2, 42, 49] provide high visual quality but require dense sampling and costly MLP inference, limiting scalability. In contrast, 3D Gaussian Splatting (3DGS) [17, 46, 6] represents scenes using explicit differentiable Gaussian primitives, achieving real-time high-quality rendering. Building on this, 2D Gaussian Splatting [14] uses planar primitives that simplify ray–primitive interaction and improve geometric stability. However, 2D and 3D Gaussian splatting represent appearance using view-dependent spherical harmonics [17, 14], without spatial color variation within each primitive, constraining its expressiveness.
@@
To address this limitation, texture-based Gaussian extensions [4, 28, 15] attach per-Gaussian spatially varying textures. However, the appearance parameterization of these methods is uniformly distributed across the texture space of each Gaussian, without adapting to local visual complexity [11, 37, 40]. This uniform allocation leads to inefficient utilization of the texture space: high-frequency regions (e.g., sharp edges, fine patterns, numerical markings) are allocated insufficient texture space, resulting in detail loss, while large smooth regions inefficiently consume capacity to represent nearly uniform color. In practice, each Gaussian can only store a low-resolution texture patch (e.g., $4\times 4\times 4$), exacerbating the expressiveness constraints imposed by uniform sampling. As shown in Fig. LABEL:fig:teaser, this results in blurred appearances in visually complex regions. Simply increasing texture resolution yields only marginal improvements at quadratically increased memory and bandwidth cost, while neural texture fields [44, 48] break the real-time rendering performance of the Gaussian-based representations.
@@
Motivated by this structural inefficiency, we propose an adaptive sampling scheme of the per-Gaussian texture where sampling should follow the local signal frequency rather than remain spatially uniform. To achieve this, we introduce a frequency-aligned differentiable texture reparameterization that continuously modulates sampling density across the texture domain. Grounded in adaptive sampling theory [12, 29, 51], this allocates more sampling capacity to high-frequency regions and less to smooth areas. Building on this formulation, we present FACT-GS, a Frequency-Aligned Complexity-Aware Texture Gaussian Splatting framework that integrates this frequency-adaptive reparameterization.
@@
Concretely, FACT-GS is built upon 2D Gaussian Splatting [14] and introduces a learnable deformation field defined in the texture domain. This field predicts continuous per-texel displacements, inducing a smooth warping whose Jacobian determinant [30] directly modulates local sampling density. In contrast to previous uniform texture parameterizations [4, 28, 15], which allocate equal resolution regardless of signal complexity, our warping adaptively concentrates samples in high-frequency regions while compressing low-frequency areas. This realizes adaptive sampling as a differentiable, per-primitive reparameterization, allowing texture capacity to naturally focus where visual detail is needed. The entire framework is trained end-to-end under standard photometric supervision, as the deformation field is designed to integrate seamlessly into the differentiable splatting pipeline. As a result, FACT-GS recovers sharper high-frequency appearance and uses parameters more efficiently, all while maintaining real-time rendering performance under the same primitive budget.
@@ 2 Related Work
Novel View Synthesis. Novel view synthesis (NVS) aims to generate photorealistic renderings of 3D scenes from unseen viewpoints, typically given multi-view images with known camera poses. Its development originates from classical SfM and MVS pipelines [33, 34, 1, 31, 32, 10, 9], which reconstruct geometry through multi-view correspondences but suffer from occlusion sensitivity and limited viewpoint extrapolation. Neural radiance fields (NeRF) [25] later revolutionized NVS by representing scenes as implicit neural functions optimized via photometric supervision. Subsequent extensions improve anti-aliasing [2], large-scale reconstruction [38, 3], reflection modeling [39], and acceleration via grid-based encodings [49], yet they remain constrained by dense ray sampling and slow MLP inference.
This motivates explicit representations such as 3D Gaussian Splatting [17], which replace neural volumes with differentiable Gaussian primitives for real-time rendering. Building upon this foundation, follow-up works improve geometric accuracy and rendering stability through enhanced kernel formulations and densification strategies [8, 20, 21]. To further refine geometry, 2D Gaussian Splatting [14] introduces planar Gaussian primitives for more accurate and efficient radiance field representation, enabling fine-grained geometric reconstruction while maintaining high appearance quality. However, both 3D and 2D Gaussian Splatting modeled appearance via view-dependent spherical harmonic coefficients, but lack spatial color variation within each primitive, limiting their expressiveness.
Spatially Varying Gaussians. Recent works introduce spatially varying Gaussians to overcome this constraint. One line of research augments each primitive with RGB or RGBA textures to model spatially varying colors [4, 43, 28, 15, 35] or surface normals for relighting [36]. Another employs neural texture fields to decouple geometry and appearance [44, 48], achieving high visual fidelity at the cost of slower training and rendering due to MLP inference.
@@
Adaptive Sampling and Frequency-Aware Parameterization. Classical rendering and Monte Carlo integration extensively explored adaptive sampling [11, 51, 29, 37, 40], where sample density is modulated by local signal frequency or reconstruction error to achieve efficient high-frequency reconstruction. At the representation level, multi-resolution texture filtering such as mipmapping and anisotropic filtering [13, 41, 27] pre-filters textures across scales to suppress aliasing and maintain visual consistency under varying viewing distances. Recent neural and Gaussian-based formulations [22, 50, 24] extend this principle to learned 3D representations, enforcing frequency- or scale-consistent parameterization in both geometry and appearance reconstruction. However, these methods operate at the global or feature-map level and still assume uniform sampling within each primitive, limiting their ability to adapt texture-space resolution to local signal variation.
@@ 3 Preliminaries
We adopt the 2D Gaussian Splatting representation [14]. Each primitive lies on a local tangent plane with center $p_{k}$, orthonormal basis $(t_{\beta},t_{\gamma},n_{k})$, and scales $(s_{\beta},s_{\gamma})$. Local coordinates on the plane are $(\beta,\gamma)$:
@@ 3.2 Textured Gaussians
Standard 2D Gaussians provide only a single view-dependent color for each primitive. Textured Gaussians [4] instead attach a learnable RGBA texture map $\mathbf{T}_{i}\in\mathbb{R}^{\tau\times\tau\times 4}$ to each primitive, enabling spatially varying appearance within tangent plane. Although originally formulated in 3D, the official implementation adopts a 2D variant for more accurate and efficient ray–Gaussian intersection, which we also follow.
@@ 4.1 Motivation
Let $C(u,v)$ denote the ground-truth color function, and let the texture $\mathbf{T}$ be its discrete representation, where texels sample $C(u,v)$ in texture space. The sampling density $\rho(u,v)$ represents the number of texels per unit area. Uniform parameterization is structurally inefficient because it assigns the same density everywhere despite large spatial variation in the underlying appearance signal [30, 51].
@@
Following prior work that uses local variation as a proxy for frequency content [5, 29], we adopt the gradient-based target density
@@ 4.2 Frequency-Aligned Complexity-Aware Texture Reparameterization
In contrast to texture-based Gaussians [4, 28], which operate on fixed uniform grids and primarily improve detail by increasing resolution, our formulation adjusts sampling density geometrically through $\Phi$ while keeping the grid resolution fixed. This enables non-uniform, complexity-aware texture parameterization that remains fully differentiable and end-to-end trainable.
@@
The local signal frequency can be approximated by the gradient magnitude $\|\nabla c_{i}^{\mathrm{tex}}(u,v)\|$ [5]. Under the warp, the uniform-grid texture gradient $\nabla\mathbf{T}_{i}^{\mathrm{RGB}}$ is transformed by the singular values of $J_{\Phi_{i}}$, which respectively amplify or attenuate its magnitude. These effects manifest as regions receiving more texels allocated or fewer texels received, enabling the FACT texture $c_{i}^{\mathrm{tex}}(u,v)$ to respond faithfully to higher-frequency variation in the ground-truth color function $C(u,v)$.
@@ 6   Sample $c_{i}^{\mathrm{tex}},\alpha_{i}^{\mathrm{tex}}$ from $\mathbf{T}_{i}(u^{\prime},v^{\prime})$;
Figure 3: Novel view synthesis comparison under reduced primitive budgets (10% / 1% of default 2DGS, exact budgets are shown in Supp. 7.3). 2DGS [14] loses high-frequency appearance, and Textured GS [4] blurs fine detail due to uniform texture sampling. Our method preserves sharp edges, textures, and material details (zoom-ins), achieving significantly higher perceptual fidelity without increasing parameters.
@@ 4.3 Optimization
Following Textured Gaussians [4], we adopt a two-stage optimization for stability. In the first stage, the 2D Gaussian primitives are trained using the standard losses from 3D Gaussian Splatting [17], combining $\mathcal{L}_{1}$ photometric and structural similarity ($\mathcal{L}_{\text{SSIM}}$) terms, augmented with a mask loss $\mathcal{L}_{\alpha}$ when ground-truth masks are available:
@@ 5.1 Implementation Details
We build on the gsplat [45] framework and extend it with a lightweight CUDA kernel for frequency-aligned texture warping. The per-Gaussian texture space sampling-density modulation described in section 4.2 introduces negligible overhead during inference. All experiments are run on a single NVIDIA RTX A6000 GPU.
@@ 5.2 Datasets and Evaluation Metrics
Dataset. We evaluate our method on five standard benchmarks for novel view synthesis: NeRF Synthetic [25] (8 scenes), MipNeRF 360 v2 [3] (7 scenes), DTU [16] (4 scenes), Tanks & Temples [18] (3 scenes), and LLFF [26] (2 scenes).
Metrics. Following Textured GS [4], we report Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS) [47] for quantitative evaluation.
@@ 5.3 Experimental Results
Methods NeRF Synthetic [25] MipNeRF 360v2 [3] DTU [16] Tanks & Temples [18] LLFF [26]
@@
Table 1: Quantitative comparison with 2DGS [14] and Textured GS [4] under varying numbers of Gaussian primitives. Our method consistently achieves higher fidelity across different primitive budgets. (Best scores are shown in bold, second-best are underlined.)
Quantitative Results. We evaluate our method against Textured GS [4] and 2D GS [14] under varying primitive budgets ($0\%, 1010\%$, $1\%$). All methods follow the same 2D GS training protocol [14]. Table 1 summarizes the results; corresponding average primitive counts are provided in the appendix. To ensure comparable parameter budgets, we match per-Gaussian texture capacity by setting $\tau_{\text{tex}}=5$ for Textured GS and $\tau_{\text{FACT}}=4$ for our method.
@@
We further evaluate our method ($100\%$ primitive budget) on the NeRF Synthetic [25] and MipNeRF 360-Indoor [3] datasets. We compare against Textured GS [4], SuperGS [43], GsTex [28], and Neural Shell [48].
As shown in Table 2, our method achieves the highest overall fidelity while maintaining real-time rendering performance (in Table 3). In contrast, methods such as Neural Shell [48] improve accuracy only by increasing computation. This indicates that our method does not rely on capacity scaling, but instead reallocates texture resolution to high-frequency regions—yielding structural efficiency gains directly predicted by adaptive sampling theory.
Detail Preservation Under Limited Primitive Budgets. Fig. 3 compares novel view synthesis when using only $10\%$ and $1\%$ of the default 2DGS primitive count. 2D GS [14] loses fine textures entirely due to its single-color appearance model, and Textured GS [4] produces blurred patterns because its uniform texture grid allocates capacity evenly across the surface. In contrast, our method preserves sharp edges, printed characters, and high-frequency materials by reallocating sampling density to regions that require detail. This confirms that frequency-aligned parameterization maintains perceptual fidelity even when representational resources are severely constrained.
@@
Methods NeRF Synthetic [25] MipNeRF 360-indoor [3]
@@
3DGS [17] 33.34 0.969 0.030 31.03 0.921 0.188
SuperGS [43] 33.71 0.970 0.031 30.23 0.917 0.188
GsTex [28] 33.37 0.965 0.041 30.46 0.915 0.204
Neural Shell [48] 33.50 0.967 0.032 30.59 0.911 0.174
Textured GS [4] (100%) 33.91 0.968 0.024 31.71 0.938 0.061
@@
Table 2: Quantitative comparison with prior Gaussian-based methods, including 3DGS [17], Neural Shell [48], SuperGS [43], GsTex [28], and Textured GS [4]. All methods are evaluated using the full (100%) 2DGS primitive budget [14] on NeRF Synthetic [25] and MipNeRF 360-Indoor [3]. (Best results are shown in bold, and second-best are underlined.)
Figure 4: Qualitative comparison under high primitive budgets ($100\%$ for Lego, $50\%$ for Church). 2D GS [14] lacks spatial texture variation, and Textured GS [4] shows blurred high-frequency patterns due to uniform texture sampling. Our method preserves thin structures and fine texture details.
Inference Speed (FPS) Textured GS [15] Neural Shell [48] Ours